# A Context-Driven Approach for Co-Auditing Smart Contracts with The Support of GPT-4 code interpreter

### v1

@@ Abstract
The surge in the adoption of smart contracts necessitates rigorous auditing to ensure their security and reliability. Manual auditing, although comprehensive, is time-consuming and heavily reliant on the auditor’s expertise. With the rise of Large Language Models (LLMs), there is growing interest in leveraging them to assist auditors in the auditing process (co-auditing). However, the effectiveness of LLMs in smart contract co-auditing is contingent upon the design of the input prompts, especially in terms of context description and code length. This paper introduces a novel context-driven prompting technique for smart contract co-auditing. Our approach employs three techniques for context scoping and augmentation, encompassing code scoping to chunk long code into self-contained code segments based on code inter-dependencies, assessment scoping to enhance context description based on the target assessment goal, thereby limiting the search space, and reporting scoping to force a specific format for the generated response. Through empirical evaluations on publicly available vulnerable contracts, our method demonstrated a detection rate of 96% for vulnerable functions, outperforming the native prompting approach, which detected only 53%. To assess the reliability of our prompting approach, manual analysis of the results was conducted by expert auditors from our partner, a world-leading smart contract auditing company [^1] The name of the company is obfuscated to respect the blind review process. The experts’ analysis indicates that, in unlabeled datasets, our proposed approach enhances the proficiency of the GPT-4 code interpreter in detecting vulnerabilities.
Index Terms:
@@
I Introduction
@@
In this work, we introduced a novel context-driven prompting technique for smart contract auditing that aims to provide appropriate wording and enough code context based on direct code dependencies to uncover vulnerabilities and provide code audit recommendations. To guide the LLM in identifying the vulnerabilities, we used the Common Weakness Enumeration (CWE)[9] to reduce the search space, in addition to a set of predefined Common Audit Questions (CAQ) extracted from previous audit reports to scope the generated reports. Moreover, to provide sufficient code context while avoiding including the entire code within the prompt, we developed a chunking approach based on code call graphs, which we refer to as Code Call List or CCL-based chunking. The method, grounded in function call graphs, presents a promising avenue for future research in LLM-based automated code reviews. For evaluation, we conduct experiments to answer the following research questions:
@@
    RQ1. How does CCL-based chunking perform in comparison to full code prompting? The objective of RQ1 is to evaluate the effectiveness of CCL-based chunking compared to full-code prompting, where GPT-4 code interpreter normally performs automated mechanical chunking by dividing the code based on its maximum token size for a prompt.
@@
    RQ2. What is the impact of different prompt phrasing approaches (CWE-based, or CAQ-based) on the effectiveness of the GPT results of CCL-based chunking?
    The objective of RQ2 is to study the impact of utilizing different wording and phrasing techniques for context augmentation based on expected results (e.g., the expected vulnerability in CWE-based) or common auditing inquiries (CAQ-based) approaches on the cost and effectiveness CCL-Chunk promoting to guide the ChatGPT to retrieve the correct results.
@@
    RQ3: To what extent does a human auditor analysis of GPT-4 code interpreter’s results on an unlabeled dataset align with our previous findings?
    The objective of RQ3 is to evaluate the reliability of our proposed prompting approach by assessing its consistency with manual auditing by expert auditors from our partner (a world-leader smart contract auditing company). A second experiment was conducted using unlabeled datasets, and expert auditors were invited to perform independent analysis and qualitative assessment of GPT-4 code interpreter’s responses, providing an additional layer of verification to our proposed methodology.
@@
    We introduce a novel context-driven prompting technique for smart contract co-auditing that aims to provide appropriate wording and enough code context based on direct code dependencies to provide code audit recommendations.
@@
    We present a dataset of vulnerable functions and their corresponding Code Call Lists, annotated by experienced auditors, facilitating empirical evaluation and future research.
@@
    We conduct comparative experiments to evaluate different prompt engineering techniques for context representation and assess the generated audits.
@@
The rest of this research paper is organized as follows: Section II provides the motivation for this works. Section III explains our methodology. Section IV presents our empirical evaluation. Section V discusses our findings. Section VI provides all the artifacts used for this work and section VII introduces related work. Section VIII outlines the threats to validity and Section IX concludes the paper.
II Smart Contract Co-Auditing
II-A Code Auditing
@@
Software development encompasses a range of software quality assurance activities, including defect management, testing and code review. These activities aim to ensure the overall quality of the software [11][12]. In the context of code review, reviewers evaluate whether the source code meets both functional and non-functional requirements. On the other hand, blockchain-based projects also requires rigorous quality assurance process to ensure the quality, security and overall reliability of the codebase. This process is called code audit. Code auditing and code review are both critical processes in software development. While they share similarities, they differ in scope, purpose, and complexity [5]. In the context of smart contracts, code auditing can be a complex process that involves a comprehensive and in-depth analysis of the whole codebase of the project with the objective of identifying vulnerabilities, potential attack vectors and security weaknesses. This process involves different tasks, including fault localization, issue explanation, documentation, and issue fix suggestions, which can be time-consuming and highly dependent on the qualifications of the auditors.
II-B Co-auditing with Generative AI
Co-auditing refers to the practice of involving both artificial intelligence (AI) systems and human oversight in the auditing process [7]. A practice that witnessed fast advancement with the emergence of LLMs. The need for co-auditing can vary on the domain, but it is crucial in applications where factual or reasoning errors can have serious repercussions [13]. An example of such an area is smart contract auditing.
@@ II-B Co-auditing with Generative AI
Recently, different studies have focused on exploring the use of LLMs for code auditing [12][10]. While these works highlight the capacity of such models for solidity code understanding and interpretation, the accuracy of LLM for the task of detecting vulnerabilities is still low. In [10], different limitations have been claimed for the use of GPT-4 code interpreter for code audits, including uncertainty and context length. The performance of GPT-4 code interpreter can degrade considerably when the prompted code becomes very long. These long contextual dependencies can lead to ambiguity in properly understanding the prompted codebase. Besides, code auditing requires knowledge about the whole code base, which makes it more complex for tools such as GPT. Although, theoretically, the standard GPT-4 code interpreter model offers 8000 tokens as a context limit, the paper argues that the maximum code size supported for the task of vulnerability detection is much lower, and the performance of GPT-4 code interpreter degrades considerably when the lines of code exceed a certain limit that is lower than the maximum token limit supported by GPT-4 code interpreter. Besides, the manual tests we performed revealed that GPT-4 code interpreter fails to detect all the vulnerabilities in a smart contract when the codebase is too long, inducing many false negatives.
@@
As explained by the authors of the paper ”What It Wants Me To Say” [14], the use of generative AI in tasks such as code generation or vulnerability detection can be challenging due to the gap between end-user programmers and code-generating LLMs, or in our case, code auditor intent and the LLM output. The authors emphasized the role of scoping, elaboration, language restructuring, and intent shaping in grounded utterance [14]. Stemming from these results, this work focuses on approaches to improving co-auditing through context scoping and augmentation. We particularly focus on the following three scopes:
@@
    Code scoping: This scoping is related to managing long code with complex inter-dependencies. In such cases, scoping should include methods to direct the LLM’s attention to the interrelated code elements that could be far apart but have close interactions. Understanding how code elements interact could lead to identifying potential vulnerabilities.
@@
    Assessment scoping: This scoping is related to defining the specific vulnerabilities, weaknesses, or security issues an auditor is targeting within the codebase. Scoping is required to limit the search space and provide better context to the LLM to trigger attention.
III Methodology
@@ III-A1 Code Flattening
Smart contract code often involves importing, calling, and referencing other contracts and libraries and implementing interfaces that extend the contract’s functionality. These dependencies can increase the complexity of auditing smart contracts, especially when dealing with multiple imported contracts. To simplify this complexity, we employ a code-flattening approach, which consolidates all the code from the contract and its dependencies into a single file. In the flattening process, we identify and include the functions within the main contract and the call relationships between them, including their interactions with interdependent functions from imported contracts. Comments are retained because they offer valuable context for understanding the intended functionality of the code. Moreover, the process ignores unreachable code for simplicity. Flattening a contract creates a single file that includes all code from the contract and its dependencies, eliminating the need to manage multiple files and ensuring that all dependencies are included correctly. To achieve this code flattening process, we utilized the Slither contract flattening tool [15], specifically employing the OneFile strategy. Slither is a static analysis tool that detects potential issues in Solidity code, including problems related to imports and interfaces. The OneFile strategy merges all contracts within the project into a single, standalone file. This approach further enhances code simplicity and readability, making it more convenient for auditing, reviewing, and maintaining the smart contract codebase. By using Slither’s OneFile strategy, we streamline the codebase and create a consolidated, easily accessible code structure for efficient analysis and development.
III-A2 Function Call Graph (FCG) Extraction

### v2

# A Context-Driven Approach for Co-Auditing Smart Contracts Using GPT-4 Code Interpreter

@@ Abstract
Context: The surge in the adoption of smart contracts necessitates rigorous auditing to ensure their security and reliability. Manual auditing, although comprehensive, is time-consuming and heavily reliant on the auditor’s expertise. With the rise of Large Language Models (LLMs), there is growing interest in leveraging them to assist auditors in the auditing process (co-auditing). However, the effectiveness of LLMs in smart contract co-auditing is contingent upon the design of the input prompts, especially in terms of context description and code length.
Objectives: This paper introduces a novel context-driven prompting technique for smart contract co-auditing. The primary objective is to improve the effectiveness of LLMs in detecting vulnerabilities in smart contracts through enhanced prompt design.
Methods: Our approach employs three techniques for context scoping and augmentation, namely: Code Scoping, which involves chunking long code into self-contained code segments based on code interdependencies; Assessment Scoping, which enhances context description based on the target assessment goal to limit the search space; and Reporting Scoping, which forces a specific format for the generated response.
Results: Through empirical evaluations on publicly available vulnerable contracts, our method demonstrated a detection rate of 96% for vulnerable functions, outperforming the native prompting approach, which detected only 53%. To assess the reliability of our prompting approach, manual analysis of the results was conducted by two expert auditors from Quantstamp, a world-leading smart contract auditing company. The experts’ analysis indicates that, in unlabeled datasets, our proposed approach enhances the proficiency of the GPT-4 code interpreter in detecting vulnerabilities.
Conclusion: The context-driven prompting technique significantly improves the effectiveness of LLMs in smart contract co-auditing. This method not only enhances the detection rate of vulnerabilities but also reduces the dependency on manual auditing, thereby making the auditing process more efficient and reliable.
keywords:
@@
journal: Information and Software Technology
[…]
[poly] organization=Department of Computer and Software Engineering, Polytechnique Montreal, city=Montreal, country=Canada
[…]
[icelandUniv] organization=Department of Computer Science, Reykjavik University, city=Reykjavík, country=Iceland
[…]
[polyhong] organization=Department of Computer Science, The Hong Kong Polytechnic University , city=Hung Hom, country=Hong Kong
[…]
[quant] organization=Quantstamp, city=Dover, country=United States
1 Introduction
@@
In this work, we introduced a novel context-driven prompting technique for smart contract auditing that aims to provide appropriate wording and enough code context based on direct code dependencies to uncover vulnerabilities and provide code audit recommendations. To guide the LLM in identifying the vulnerabilities, we used the Common Weakness Enumeration (CWE)[6] to reduce the search space, in addition to a set of predefined Common Audit Questions (CAQ) extracted from previous audit reports to scope the generated reports. Moreover, to provide sufficient code context while avoiding including the entire code within the prompt, we developed a chunking approach based on code call graphs, which we refer to as Code Call List or CCL-based chunking. For evaluation, we conduct experiments to answer the following research questions:
@@
    RQ1. How does CCL-based chunking perform in comparison to full code prompting? The concept of chunking, particularly CCL-based chunking, is inspired by the manual auditing process employed by expert auditors. During the audit of a smart contract, auditors meticulously verify each function to ensure it is free from vulnerabilities. This necessitates a comprehensive understanding of all the code executed within each code snippet. For Research Question 1 (RQ1), we assess the performance of CCL-based chunking relative to full-code prompting. In full-code prompting, the GPT-4 code interpreter typically executes automated mechanical chunking by dividing the code based on the maximum token size permissible for a single prompt. Our evaluation aims to determine the effectiveness of CCL-based chunking in comparison to this automated approach.
@@
    RQ2. What is the impact of different prompt phrasing approaches (CWE-based, or CAQ-based) on the effectiveness of the GPT results of CCL-based chunking? The objective of RQ2 is to study the impact of utilizing different wording and phrasing techniques for context augmentation based on expected results (e.g., the expected vulnerability in CWE-based) or generic auditing inquiries (searching for all vulnerabilities in CAQ-based) approaches on the cost and effectiveness of CCL-based chunking to guide the ChatGPT to retrieve the correct results.
@@
    RQ3: To what extent does a human auditor’s analysis of GPT-4 code interpreter’s results on an unlabeled dataset align with our previous findings? The objective of RQ3 is to evaluate the reliability of our proposed prompting approach by assessing its consistency with manual auditing by expert auditors from our partner (a world-leading smart contract auditing company). A second experiment was conducted using unlabeled datasets, and expert auditors were invited to perform an independent analysis and qualitative assessment of GPT-4 code interpreter’s responses, providing an additional layer of verification to our proposed methodology.
@@
    We introduce a context-driven approach to enhance LLM performance in smart contract auditing by proposing three scoping techniques: Code Scoping, Assessment Scoping, and Reporting Scoping. These techniques aim to improve understanding and generate accurate audit reports by addressing the challenges of (i) long code contexts in prompts, (ii) context related to the audit process, and (iii) context related to expected target responses.
@@
    For each scoping technique, we provide a practical implementation. Code Scoping is realized through CCL (Code Call List), which organizes code into logical chunks based on function call dependencies to improve contextual awareness. Assessment Scoping is realized through CAQ (Common Audit Questions), a tailored set of questions that guide the LLM’s focus on specific vulnerabilities. Reporting Scoping is realized through CWE (Common Weakness Enumeration), helping the LLM categorize and report vulnerabilities using standardized industry classifications.
@@
    We present hypotheses regarding the effectiveness of the scoping techniques and empirically test these through experiments to assess their impact on audit quality.
  4. 4.
    We provide a new dataset of vulnerable functions, manually labeled by two professional smart contract auditors, to aid in the evaluation and validation of the proposed methodologies.
@@
The rest of this research paper is organized as follows: Section II provides the motivation for this work. Section III explains our methodology. Section IV presents our empirical evaluation. Section V discusses our findings. Section VI provides all the artifacts used for this work and section VII introduces related work. Section VIII outlines the threats to validity and Section IX concludes the paper.
2 Smart Contract Co-Auditing
2.1 Code Auditing
@@
Software development encompasses a range of software quality assurance activities, including defect management, testing, and code review [10, 14]. In the context of code review, reviewers evaluate whether the source code meets both functional and non-functional requirements. On the other hand, blockchain-based projects also require a rigorous quality assurance process to ensure the quality, security, and overall reliability of the codebase. This process is called code audit. Code auditing and code review are both critical processes in software development. While they share similarities, they differ in scope, purpose, and complexity [13]. In the context of smart contracts, code auditing can be a complex process that involves a comprehensive and in-depth analysis of the whole codebase of the project with the objective of identifying vulnerabilities, potential attack vectors, and security weaknesses. This process involves different tasks, including fault localization, issue explanation, documentation, and issue fix suggestions, which can be time-consuming and highly dependent on the qualifications of the auditors.
2.2 Co-auditing with Generative AI
@@
Co-auditing refers to the practice of involving multiple auditors, which can include both artificial intelligence (AI) systems and human oversight, in the auditing process. [39]. A practice that witnessed fast advancement with the emergence of LLMs. The need for co-auditing can vary depending on the domain, but it is crucial in applications where factual errors can have serious repercussions [15]. An example of such an area is smart contract auditing. Recently, different studies have focused on exploring the use of LLMs for code auditing [14, 3]. While these works highlight the capacity of such models for Solidity code understanding and interpretation, the accuracy of LLMs for the task of detecting vulnerabilities is still low. In [3], different limitations have been reported for the use of GPT-4 code interpreter for code audits, including uncertainty and context length. The performance of GPT-4 code interpreter can degrade when the prompted code becomes very long. Long contextual dependencies can lead to ambiguity in properly understanding the prompted codebase. Besides, code auditing requires knowledge about the whole code base, which makes it more complex for tools such as GPT.
Despite the theoretical capacity of large language models (LLMs) like GPT-4 to process up to 8000 tokens, our research reveals that their efficacy in smart contract vulnerability detection is hindered by limitations in contextual understanding, particularly when encountering complex code structures. Notably, the performance degradation of LLMs in this domain is not solely attributable to context length, but rather a multifaceted issue encompassing the model’s ability to capture nuanced code relationships, reason about indirect dependencies, and generalize to novel contract architectures. The manual tests we performed revealed that GPT-4 code interpreter fails to detect all the vulnerabilities in a smart contract when the codebase is too long, inducing many false negatives. This underscores the need for innovative prompting techniques that can effectively harness the strengths of LLMs while mitigating their limitations in smart contract auditing. The integration of generative AI in code auditing poses a unique challenge, as highlighted by Liu et al. in ”What It Wants Me To Say” [20]. The disconnect between the end-user programmer’s requirements and the output of large language models (LLMs) can lead to suboptimal results in tasks like code generation and vulnerability detection. In the context of code auditing, this gap can have significant implications, as LLMs may not fully comprehend the nuanced security concerns and auditing standards that human auditors take for granted. However, LLMs can still play a vital role in augmenting the code auditing process, particularly in tasks such as:
  • 1.
    Identifying potential vulnerabilities and anomalies in large codebases
  • 2.
    Providing recommendations for code improvements and security enhancements
@@
The authors emphasized the role of scoping, elaboration, language restructuring, and intent shaping in grounded utterance [20]. Stemming from these results, this work focuses on approaches to improving co-auditing through context scoping and augmentation. We introduce the following three scopes:
@@
    Code scoping: This scoping is related to managing long code with complex inter-dependencies. In such cases, scoping should include methods to direct the LLM’s attention to the interrelated code elements that could be far apart but have close interactions. Understanding how code elements interact could lead to identifying potential vulnerabilities. While we do not explicitly quantify this relationship in the paper, the concept involves recognizing and structuring the context based on inter-dependencies, which can be informed by function call graphs and dependency analysis as explained in Section 3.1.
@@
    Assessment scoping: This scoping is related to defining the specific vulnerabilities, weaknesses, or security issues an auditor is targeting within the codebase. For example, the auditor may target issues like reentrancy, integer overflows, access control vulnerabilities, or gas limit issues. This focus helps narrow down the scope of the context. This scoping technique is closely related to retrieval-augmented approaches, as it helps guide the LLM to focus on the most relevant information by retrieving context-specific details for accurate vulnerability detection. While we do not explicitly define formal parameters or metrics for Assessment Scoping at this stage, in the implementation of this scoping technique in section xx the context was identified based on targeted vulnerabilities retrieved from the Common Weakness Enumeration (CWE) database and evaluated by accuracy of vulnerability detection
Researchers have been investigating various aspects of context in large language models (LLMs), including how it is defined, utilized, and evaluated [19, 17]. In [19], authors address the challenge of long-context understanding for LLMs. The paper defines the context window as the maximum amount of text or tokens that a model can process in a single input. The study revealed that LLMs excel in short dependency tasks but struggle with more complex long dependency tasks, indicating a gap in true long-context understanding. Yilun Zhu et al. [41] conducted an empirical investigation into the contextual comprehension capabilities of large language models (LLMs). Their systematic evaluation assessed the ability of LLMs to accurately interpret subtle and implicit contextual information, a crucial factor in tasks requiring profound contextual awareness. The findings revealed that while LLMs exhibit satisfactory performance in general linguistic tasks, they frequently misinterpret or fail to comprehend nuanced contextual cues. This study underscores the significance of contextual definition and structure in optimizing LLM performance, particularly for code generation and vulnerability detection tasks. Notably, LLM performance degrades when processing large codebases, characterized by intricate dependencies and numerous function calls, even when token counts remain within contextual windows [18]. For comprehensive codebase understanding, semantic relationships between disparate code segments assume paramount importance in structuring input prompts. Harnessing the structural properties of function call graphs, throw code scoping, offers a promising approach to represent logical connections between functions, thereby enhancing contextual awareness.
To inform the development of effective Large Language Model (LLM) auditing queries for smart contracts, we employed a mixed-methods approach combining card sorting methodology and expert elicitation. Our investigation involved analyzing a dataset of audit reports extracted from Code4Arena, a prominent capture-the-flag platform utilized by the auditing community, to identify key themes, patterns, and vulnerabilities prevalent in smart contract audits. In collaboration with professional auditors, we distilled these findings into a comprehensive set of tailored questions designed to elicit relevant information from the LLM, ensuring alignment with industry auditing standards. By integrating human expertise and real-world audit report data, our approach establishes a robust foundation for exploring the potential of LLMs in automated smart contract auditing, providing a systematic and informed basis for enhancing the efficiency and effectiveness of this critical process.
This study integrates Large Language Models (LLMs) into real-world smart contract auditing workflows, enhancing auditing practice by tailoring LLM outputs to specific vulnerability detection and reporting requirements. By bridging the gap between LLM general capabilities and smart contract security needs, our approach addresses specifically auditing smart contracts. Our focused methodology provides actionable insights into LLM applicability, rather than broadly investigating LLM behavior.
In the evaluation section, we provide an example of implementation of our understanding of the context definition for the task of auditing smart contract.
2.3 Motivation Example
Listing 1 illustrates an example of a smart contract called TokenizingVault. This vault contract takes users’ diverse ERC20s (token standard) deposits and then issues corresponding ”credits” as deposit tokens. These tokens can then be exchanged to reclaim the initial deposit. This design provides fungibility that allows users to trade their secured assets seamlessly. The TokenizingVault contract imports two other contracts, namely ERC20.sol and ERC20CreditToken.sol. ERC20 and ERC20CreditToken also import different other contracts which we tried to prompt this contract along with the other imported contract. The contracts source code length exceeds 8000 token supported by GPT4. A direct solution for such a challenge is a naive truncation. We prompted the source code of the contracts sequentially in different prompts. We used a binary classification of the contract and asked GPT to locate the vulnerable functions, if any, and explain the vulnerability.
Listing 1: TokenizingVault contract
pragma solidity 0.8.15;
import "@openzeppelin/contracts/token/ERC20/ERC20.sol";
import {ERC20CreditToken} from "./ERC20CreditToken.sol";
contract TokenizingVault {
/// ERC20 accounting tokens (unique per underlying)
mapping(ERC20 => ERC20CreditToken) public creditTokens;
/// ERC20CreditToken reference implementation
ERC20CreditToken public immutable creditTokenImpl;
constructor() {
creditTokenImpl = new ERC20CreditToken();
function create(ERC20 underlying, uint256 amount)
external nonReentrant returns (ERC20CreditToken, uint256)
ERC20CreditToken creditToken = creditTokens[underlying];
// Revert if no token exists, must call deploy first
if (creditToken == ERC20CreditToken(address(0x00)))
revert('Token Does Not Exist');
// Transfer in underlying
underlying.transferFrom(msg.sender, address(this), amount);
// Mint new credit tokens
creditToken.mint(msg.sender, amount);
return (creditToken, amount);
function redeem(ERC20CreditToken token, uint256 amount)
external nonReentrant
token.burn(msg.sender, amount);
token.underlying().transfer(msg.sender, amount);
function deploy(ERC20 underlying)
external nonReentrant returns (ERC20CreditToken)
// Create credit token if one doesn't already exist
ERC20CreditToken creditToken = creditTokens[underlying];
if (creditToken == ERC20CreditToken(address(0))) {
bytes memory tokenData = abi.encodePacked(underlying, address(this));
creditToken = ERC20CreditToken(address(creditTokenImpl).clone(tokenData));
creditTokens[underlying] = creditToken;
return creditToken;
Due to the complex dependency, the code base is prompted using a sequence of 5 prompts. GPT is composed of 3 parts: general observations, Potential Vulnerabilities and Suggestions, and conclusions. As we can see in the generated response [^1] https://chat.openai.com/share/fef0be4f-0def-449a-bf22-30fd9cc373cb, GPT fails to detect a crucial vulnerability which is related to the redeem function.
The core issue for this contract is the function redeem. The function redeem is external and does not have an access control list or input validation schema. This function is vulnerable because it does not validate if the passed token is authenticated. One exploitation scenario is that an attacker could invoke the redeem function using a maliciously crafted ERC20CreditToken. If this token’s underlying method points to a legitimate and valuable ERC20 token that other users have stored in the vault, it may lead to unauthorized withdrawals. The vulnerable function is only three lines long. Although GPT’s response mentioned different code issues related to the contract, the vulnerability has not been listed. We repeated the experiment various times. The potential reasons can be:
  • 1.
    Lack of context depth: In the absence of specific context description, llm is capable of generating outputs in formats that are not uniformly structured. This irregularity in the generated response presents substantial obstacles in extraction of valuable insights that are both accurate and coherent. The absence of the precise description of the desired output in term of the context inquiry process, related to the questions that needs to be answered, and in term of the generated response, related to expected results template, can lead GPT to generate high level results that can increase the rate of false positive/negative and/or missing certain complex vulnerabilities.
  • 2.
    Sensitivity to source code length: The length of the prompt context and the code can affect the performance of LLMs for vulnerability detection [6, 3], where the effective maximum token number for GPT-4 is lower than 8000.
This suggests that a more effective chunking technique for the source code is required to fully exploit the LLMs code understanding capabilities. Hence, this paper proposes a new chunking technique based on a function call graph for source code. Moreover, the paper will present comparative experiments involving different groups of prompt engineering techniques for context representation and assess the generated audits, demonstrating the current advantages and limitations of the context length on the performance of LLMs for smart contract audits.
3 Implementation of Context-Driven Auditing Method
@@ 3.1.1 Code Flattening
Smart contract code often involves importing, calling, and referencing other contracts and libraries and implementing interfaces that extend the contract’s functionality. These dependencies can increase the complexity of auditing smart contracts, especially when dealing with multiple imported contracts. To simplify this complexity, we employ a code-flattening approach, which consolidates all the code from the contract and its dependencies into a single file. The flattening process consist of combining all the solidity files, dependencies , imported libraries, and merging them into one continuous file. This include parsing the solidity files, resolve imports and dependencies to make sure that all the code in included in the flattened file. In the flattening process, we identify and include the functions within the main contract and the call relationships between them, including their interactions with interdependent functions from imported contracts. Comments are retained because they offer valuable context for understanding the intended functionality of the code. Moreover, the process ignores unreachable code for simplicity. Flattening a contract creates a single file that includes all code from the contract and its dependencies, eliminating the need to manage multiple files and ensuring that all dependencies are included correctly. To achieve this code flattening process, we utilized the Slither contract flattening tool [11], specifically employing the OneFile strategy. Slither is a static analysis tool that detects potential issues in Solidity code, including problems related to imports and interfaces. The OneFile strategy merges all contracts within the project into a single, standalone file. By using Slither’s OneFile strategy, we streamline the codebase and create a consolidated, easily accessible code structure for efficient analysis and development.
3.1.2 Function Call Graph (FCG) Extraction